Given the intensity of an earthquake wave traveling through the Earth and detected 200 km from the source is 2 MW/m^2. What is the intensity of that wave if detected 300 km from the source?


1
Expert's answer
2021-07-19T09:52:25-0400

The intensity of the earthquake wave is inversely proportional to the square of distance from the source, hence

"I_2=I_1\frac{r_1^2}{r_2^2}\\
=2.0\:\rm MW/m^2\frac{(200\:\rm km)^2}{(300\:\rm km)^2}=0.89\:\rm MW/m^2"